OpenStudy (mayankdevnani):

Logical Question :

OpenStudy (mayankdevnani):

\[5060.50\div45^2=?-7.5\]

OpenStudy (lgbasallote):

how is that logic?

OpenStudy (unklerhaukus):

\[?=5060.50\div 45^2+7.5\]

OpenStudy (mayankdevnani):

it is a reasoning question

OpenStudy (mayankdevnani):

yaa!! then

OpenStudy (anonymous):

119.9556 :)

OpenStudy (unklerhaukus):

almost ten

OpenStudy (mayankdevnani):

yaa!! correct

OpenStudy (mayankdevnani):

howz that.... @UnkleRhaukus

OpenStudy (lgbasallote):

kindly point me to the reason i'm missing.....

OpenStudy (anonymous):

logic in math is abstract and discrete its a field of study in its own,

OpenStudy (mayankdevnani):

sorry...

OpenStudy (mayankdevnani):

any shortcut in this

OpenStudy (mayankdevnani):

@UnkleRhaukus how do you do this?

OpenStudy (mayankdevnani):

@medkl it was wrong!!!

OpenStudy (anonymous):

just take ? as x,then solve traditionally

OpenStudy (anonymous):

How did you get that answer @medkl ?? I am also getting the same @UnkleRhaukus

OpenStudy (unklerhaukus):

\[?=5060.50\div 2025+7.5\]\[?=\frac{5060.50}{2025}+7.5\]\[?=5060.50\div 2025+7.5\]\[?=\frac{2\times2025+1010.5}{2025}+7.5\]\[?=2+\frac{1010.5}{2025}+7.5\]

OpenStudy (mayankdevnani):

@UnkleRhaukus in 4 line i don't understand......

OpenStudy (unklerhaukus):

\[2\times2025=4050\] \[2\times2025+1010.5=4050+1010.5=5060.50\]

OpenStudy (mayankdevnani):

but how can you know that?

OpenStudy (mayankdevnani):

any special property apply here.....

OpenStudy (mayankdevnani):

@UnkleRhaukus

OpenStudy (mayankdevnani):

then i understand

OpenStudy (unklerhaukus):

i dunno

OpenStudy (mayankdevnani):

then @UnkleRhaukus

OpenStudy (mayankdevnani):

anyway... thnx

OpenStudy (anonymous):

I think it is just : Dividend = Quotient x Divisor + Remainder : 5060.5 = 2 x 2025 + 1010.5 That is why @UnkleRhaukus has written 5060.5 as 2 x 2025 + 1010.5 in the numerator of 4th line...

OpenStudy (mayankdevnani):

thnx.. @waterineyes

OpenStudy (anonymous):

You are welcome... @mayankdevnani